OpenStudy (anonymous):

Can someone explain this problem to me as you do it. I have many more and I need to understand how to do it. Find the fixed points of a function. f(x)=x^2-4x+6

OpenStudy (anonymous):

\[x ^{2}-4x+6=x\] \[x ^{2}-5x+6=0\] \[(x-2)(x-3)\] =\[x ^{2}-5+6\] x=2 x=3 Points (2,2) (3,3)

OpenStudy (anonymous):

I don't know if this is right but I hope so